OpenStudy (anonymous):

What is c over c to the 3rd power. It's a fraction

OpenStudy (anonymous):

\[\frac{c}{c^3}=\frac{1}{c^{3-1}}=\frac{1}{c^2}\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw